OpenStudy (anonymous):

Manually calculate the compound interest on an investment of $8,000 at 6% interest, compounded semiannually, for 18 months. a.$783.54 b.$788.18 c.$8,759.35 d.$8,788.18

OpenStudy (anonymous):

divide the interest rate by 72 .6 divided by 72

OpenStudy (anonymous):

Is that all I need to do it

OpenStudy (anonymous):

.6/72=.0083333333

OpenStudy (ranga):

Annual interest rate = 6% = 0.06. Compounded semi-annually (meaning every 6 months) Semi-annual interest rate = 0.06 / 2 = 0.03 Interest after first 6 months = p*r*t = 8000 * 0.03 * 1 = 240 Total amount after 6 months = principal + interest = 8000 + 240 = 8240 Interest after another 6 months: P*r*t = 8240 * 0.03 * 1 = ? add the interest to 8240 to get the amount. do the same calculations one more time for another 6 months to get a total of 18 months.

OpenStudy (ranga):

They are interested only in the total interest and not the total amount. So in the end add up just the interest for the 3 time periods.

OpenStudy (anonymous):

look up the principle of 72 in ref of marketing

OpenStudy (anonymous):

I got 8,487.2 is this right

OpenStudy (anonymous):

yes i belive so

OpenStudy (anonymous):

that isn't one of the answers

OpenStudy (anonymous):

Ok thanks is that I need to do

OpenStudy (anonymous):

Is that all I need to do

OpenStudy (anonymous):

I got (8000)(1+0.03)^3 - 8000 = 741.82, which should be right but isn't listed. Three six-month periods at 3% compounded.